An object is thrown vertically upward from a height of h0 ft with an initial speed of v0 ft/sec. Its height h (in feet) after t seconds is given by

h= -16t² + v0t + h0. Given this, if it is thrown vertically upward from the ground with an initial speed of 64 ft/sec,

(a) At what time will the height of the ball be 15 ft? (two answers)

(b) How long will it take for the ball to reach 63 ft?


1
Expert's answer
2021-04-15T07:19:59-0400

(a)"15=-16t^{2}+64t"


"=16t^{2}-64t+15"


"t= \frac {-b± \sqrt{b^2-4ac}}{2a}"


"t=\frac{64±\sqrt{4096-960}}{2\times16}"


"t=\frac{64+56}{32}=3.75"


"t=\frac{64-56}{32}=0.25"


3.75 sec and 0.25 sec


(b)"63=-16tt^{2}+64t"


"16t^{2}-64t+63"


"t=\frac{64± \sqrt{4096-4032}}{2\times16}"


"t=\frac{64+8 }{32}=2.25"


"t=\frac{64-8}{32}=1.75"


2.25 sec

1.75 sec